A train travelling along a straight track starts from rest at point

A and accelerates uniformly to 20 m s–1 in 20 s. It travels at

this speed for 60 s, then slows down uniformly to rest in 40 s

at point C. It stays at rest at C for 30 s, then reverses direction,

accelerating uniformly to 10 m s–1 in 10 s. It travels at this speed

for 30 s, then slows down uniformly to rest in 10 s when it

reaches point B.

a Plot a graph of the motion of the train.

b Use your graph to calculate:

i the train’s displacement from point A when it reaches

point C

ii the train’s displacement from point A when it reaches

point B

iii the train’s acceleration each time its speed changes.


1
Expert's answer
2021-03-29T08:58:21-0400

a)



b) i)


"s_1=0.5(60+120)(20)=1800\ m\\"

ii)


"s_2=1800-0.5(30+50)(20)=1400\ m"

iii) From 0 s to 20 s


"a=\frac{20-0}{20}=1\frac{m}{s^2}"

From 80 s to 120 s


"a=\frac{0-20}{40}=-0.5\frac{m}{s^2}"


From 150 s to 160 s


"a=\frac{-10-0}{10}=-1\frac{m}{s^2}"


From 190 s to 200 s


"a=\frac{20-(-10)}{10}=1\frac{m}{s^2}"